### v2

# Emerging Platforms Meet Emerging LLMs: A Year-Long Journey of Top-Down Development

Abstract
Deploying machine learning (ML) on diverse computing platforms is crucial to accelerate and broaden their applications. However, it presents significant software engineering challenges due to the fast evolution of models, especially the recent Large Language Models (LLMs), and the emergence of new computing platforms. Current ML frameworks are primarily engineered for CPU and CUDA platforms, leaving a big gap in enabling emerging ones like Metal, Vulkan, and WebGPU.
While a traditional bottom-up development pipeline fails to close the gap timely, we introduce TapML, a top-down approach and tooling designed to streamline the deployment of ML systems on diverse platforms, optimized for developer productivity. Unlike traditional bottom-up methods, which involve extensive manual testing and debugging, TapML automates unit testing through test carving and adopts a migration-based strategy for gradually offloading model computations from mature source platforms to emerging target platforms. By leveraging realistic inputs and remote connections for gradual target offloading, TapML accelerates the validation and minimizes debugging scopes, significantly optimizing development efforts.
TapML was developed and applied through a year-long, real-world effort that successfully deployed significant emerging models and platforms. Through serious deployments of 82 emerging models in 17 distinct architectures across 5 emerging platforms, we showcase the effectiveness of TapML in enhancing developer productivity while ensuring model reliability and efficiency. Furthermore, we summarize comprehensive case studies from our real-world development, offering best practices for developing emerging ML systems.
Index Terms:
Developer Productivity, Software Testing, Machine Learning Systems
$\bigstar$\$$\bigstar$\$footnotetext: Equal contribution.
I Introduction
Machine learning (ML) has been ever-evolving to advance the performance of various data-driven applications, from the early days of simple linear classifiers to the latest advancements in large language models [1]. To power large-scale ML, the silicon industry has also been continually expanding the boundaries of what is computationally possible, by innovating the computing platforms for CPUs, GPUs, and emerging domain-specific accelerators [2, 3].
Because there is a huge gap in mapping high-level ML mathematics to low-level machine executables, substantial efforts have been made to develop ML optimizers [4, 5, 6] to efficiently run ML models. These frameworks, given a trained model, deploy it over inference-optimized platforms by providing efficient native kernels (e.g., CUDA/C functions) to compute the operators/sub-graphs of the model. These kernel functions can be either manual-crafted in a native programming language (e.g., CUDA) or compiler-generated from intermediate representations [7, 8, 9, 10] (IRs). Yet, existing ML frameworks have been primarily implemented and optimized for the CPU and CUDA platforms, leaving a blank space for other emerging platforms.
With recent advances and interest in Large Language Models (LLMs), powerful new models of various architectures are released rapidly by large organizations (Figure 1). Meanwhile, an array of emerging computing platforms have been proposed recently to unlock the possibility of running ML models on environments without NVIDIA GPUs, such as Metal for Apple Silicon [11], WebGPU for web browsers [12], etc. It is crucial to enable sustainability by enabling a broader set of emerging platforms both locally and on the server. Emerging local ML provides a fundamental guarantee for user privacy and resolves the cost and network requirements of a server platform. Meanwhile, expanding server-grade computing options for GPU customers offers greater diversity in terms of performance, cost-effectiveness, flexibility, etc.
Engineering an ML framework is hard, and engineering an ML framework to enable emerging ML models on emerging platforms is an even more challenging software engineering problem. The immediate but unavoidable challenge is to learn the new programming primitives from the target platform. Furthermore, software developers can spend 35-50% of their time testing and debugging software [13]. This is also naturally true here as two main challenges arise for testing and debugging. First, sufficient and diverse test cases are required to test the implementation of individual operators, yet oftentimes developers have to craft these tests manually and the whole process is tedious and laborious. Second, the tooling support and ecosystems for emerging platforms are extremely limited (e.g., no debuggers), leading to a painful debugging and analysis experience.
Figure 1: Timeline of selected new model supports in TapML [^2] For clarify we only visualize a subset of important models we supported..
Figure 2: Bottom-up v.s. Top-down approaches to emerging ML system development
Bottom-up ML deployment. Existing ML frameworks have been commonly practicing a bottom-up development style, which unfortunately no longer meets the productivity requirement when supporting emerging ML systems. The top half in Figure 2 illustrates the key steps for bottom-up ML deployment. There are two steps to support a new model in a bottom-up fashion: (i) Developers implement previously unsupported operators; and (ii) implemented operators are put together into a whole model for end-to-end testing and debugging. Specifically, in step one, developers analyze operators to support, implement them, and manually craft tests for validation. Because manually crafted tests often use synthetic inputs whose distribution can substantially differ from those inherited from realistic inputs, operator-wise validation is inadequate. End-to-end model-wise testing is still effective in manifesting obscure bugs such as precision errors; however, bugs detected by model-wise testing are hard to debug, as the symptoms are simple pass-or-fail signals in terms of result consistency, missing details to infer the error-inducing components. As such, troubleshooting the large model space leads to challenging debugging experiences and significantly slows down the development process.
Top-down ML deployment. To this end, we propose TapML, a top-down approach and tooling to enable emerging ML systems, optimized for developer productivity. The bottom half of Figure 2 illustrates the workflow of TapML. At the high level, TapML regards the deployment as a migration task to gradually offload the model implementation from a mature source platform (e.g., CUDA) to an emerging target platform (e.g., WebGPU). Specifically, TapML takes the following steps to deploy an ML model for an emerging platform. First, TapML automates the curation of unit tests through test carving [14], which automatically tracks intermediate computation snapshots of the source-platform execution as test cases for the target-platform implementation. Notably, the inputs to bootstrap the model during instrumentation are realistic inputs (e.g., meaningful sentences for LLMs) such that the precision regression can focus on real-world data distribution. Next, we perform gradual target offloading. We start with a compute graph fully implemented on the source platform which can correctly pass all system tests. Then we gradually offload each operator in the compute graph from the source platform to the target platform, until the target platform fully runs the compute graph. For each operator, we first use the carved unit tests to perform operator-wise validation, after which we substitute the source-platform implementation with the target-platform implementation. Such a flow also provides an error isolation mechanism where a model-wise failure must be introduced by the operator newly implemented on the target platform. As such, developers can quickly localize the error in a small scope and use immediate operator- and model-wise validation feedback for efficient debugging. Meanwhile, to serve the hybrid graph and perform differential testing [15], TapML builds a universal runtime that unifies cross-platform data communication and function calling. The runtime allows us to invoke remote functions on the device and transfer its inputs and outputs between source and target platforms seamlessly in Python as if running a local program. Lastly, the workflow finalizes when the compute graph is fully implemented on the target platform and passes model-wise validation. As is shown in Figure 1, we have been applying TapML starting from 2024 Q1 and have productively supported various emerging ML models on major emerging platforms, unlocking powerful LLM access to users of various devices, even within the same day of model release.
This paper summarizes our lessons in building and deploying real-world open-source ML systems for emerging platforms. Below we summarize our contributions:
  • Dimension: While the optimization and abstraction for ML computation have been well studied in the past decade, we look at the productivity challenge for deploying emerging ML and address this challenge by advancing the software development methodology.
  • Methodology: We propose TapML, a top-down approach to deploying emerging ML on emerging platforms. While the traditional bottom-up scheme suffers from coarse-grained troubleshooting, TapML automates unit testing by carving tests from executions in mature platforms and adopts a migration-based approach to gradually offload the computation from the source platform to the target platform via our universal runtime.
  • Study: We turn our year-long experience of deploying 82 emerging models on 5 novel platforms into summarized case studies, to facilitate the future development of emerging ML systems.
II Challenges
In this section, we demonstrate key challenges and pain points in developing emerging ML systems.
II-A Insufficient and Laborious Testing
Figure 3: Sample unit test in TensorFlow.js.
Developers invest significant effort in writing extensive test suites to ensure the reliability of machine learning (ML) frameworks. For example, 36% source code of PyTorch [5], a well-known ML framework, and 42% source code of TVM [6], a well-known ML compiler, are testing code, as of Q1 2023. As an instance of emerging platforms, TensorFlow.js [16] supports 85 common operators for the WebGPU platform using 18,001 lines of source code. Despite that not all operators but a few important ones (e.g., matmul and conv2d) are unit-tested, 3,521 lines of the source code are making unit tests.
Figure 3 exemplifies a test case in TensorFlow.js which validates the functionality of conv2d. The test case is crafted by developers by specifying operator attributes (Line 2-3), the input data (Line 4-7), and the oracle including the expected output shape (Line 11) and output values (Line 12-14). Yet, this test only covers one path of the operator whereas tons of other tests use different input and operator configurations to exercise different paths. Meanwhile, the manually specified inputs are rather synthetic given the apparent value orders in x and w, which may not accurately reflect realistic data distribution for a specific model, introducing a distribution gap between test cases and real-world scenarios.
This example highlights two pain points in test development: (i) the labor-intensive nature of manual testing; and (ii) the gap between synthetic and realistic inputs. In addition to the number of tests desired, crafting comprehensive test suites requires meticulous consideration of detailed input-output shapes and values, often involving complex operator constraints [17, 18]. Furthermore, relying solely on synthetic inputs for operator validation may not accurately assess the performance of operators with real-world data, especially concerning factors like precision. This gap is particularly concerning when applying quantization techniques [19, 20] to optimize model efficiency at the cost of robustness [21].
To conclude, addressing the challenges in test development for emerging ML systems is crucial to enhance developer productivity and ensure model reliability. Efforts to bridge the gap between synthetic and realistic inputs, alongside automating test development, are essential for the continued advancement and adoption of ML models and platforms.
Figure 4: The computation representations in the TapML workflow. To deploy a trained model, TapML converts it into a general compute graph, where each operator can be offloaded to various backends through target code generation.
II-B Obscure End-to-End Bugs
Unit tests do not universally ensure the accuracy of end-to-end models due to the inability to guarantee numerical precision for every possible input. This limitation is exemplified by an observed discrepancy on Qualcomm Hexagon backends: when performing the calculation $1035-1031$ with the float16 data type, the expected result of 4 was erroneously computed as 5. Although this deviation, less than 0.1%, may seem reasonable, it diverges from the behaviors observed on other computational platforms and can have significant implications for the integrity of end-to-end models. It is important to note that such errors are input-specific and not attributable to faults in the kernel implementations themselves, rendering them elusive to detection via operator-level testing.
Furthermore, as ML models increase in complexity and size, incorporating a greater number of operators, the cumulative impact of precision errors becomes more pronounced. Even if an individual operator’s error falls within the tolerable margin of 0.1% at the operator level, the aggregate effect across hundreds of operators can amplify this error to 1% or even 5%. Such levels of inaccuracy are untenable, highlighting the need for rigorous testing protocols that account for the variances in floating-point computations attributable to different hardware and software configurations.
II-C Limited Tooling in Arising Systems
The infrastructure stack in mature platforms has been well-engineered which enables developers to easily invoke device functions, print remote data, transfer data between host and device, and so on. Moreover, mature platforms even develop advanced debugging utilities. For example, Polygraphy [22] is a graph reducer in TensorRT [4]. When a model fails to compile by TensorRT, Polygraphy automatically reduces the full model to a small sub-graph that can reproduce the error, helping developers to quickly localize the fault.
In contrast, because emerging platforms are newly established, the tooling in their ecosystem is rather limited. Consequently, oftentimes it is challenging for developers to reason about internal states of computation performed in such arising platforms, which is important for debugging. Such a fact also deteriorates the prior two challenges, solving which requires advanced tooling to communicate with the underlying platform. Therefore, it is crucial to build ideally a universal middle-end tooling that allows developers to easily inspect and analyze the model execution.
III Design of TapML
In this section, we detail the design of TapML.
III-A Overview
The goal of deploying a model on a specific platform is to describe the model’s computation using the native program of the target platform. Specifically, Figure 4 illustrates a typical deployment workflow in TapML. The input to the deployment flow is a trained model produced by the training framework (e.g., PyTorch). Subsequently, the trained model is translated by the framework into an internal intermediate representation (IR) that generally describes the semantics of the model. Through a series of general-purpose optimizations, the IR flows to the target code generation stage which transforms the computation to native programs of the target platform. For example, existing frameworks often specialize the code generation pipeline to CUDA, a mature platform in the NVIDIA ecosystem, and these computations are finally represented by CUDA C functions, shown in the purple box in Figure 4.
We describe the goal and overall steps of TapML using the example of Figure 4. The goal of TapML is to facilitate the process of implementing and testing target code generation of the model operators on the emerging platform, i.e., on WebGPU such that users can run their model on a browser, as is shown in the blue box in Figure 4. To make this happen, in Section III-B TapML makes basic assumptions for the deployment framework, including a compute graph abstraction and a reference implementation from a mature platform, i.e., CUDA here. To make testing easy, in Section III-C, TapML automatically creates realistic operator-level tests through test carving, which instruments the execution of each operator on CUDA and collects its input-output pairs as unit tests for the WebGPU support. Specifically, TapML views the deployment process as a migration task — in the very beginning we have a reliable CUDA implementation (i.e., pure CUDA compute graph) of the model and we want to eventually run the model fully on WebGPU (i.e., pure WebGPU compute graph). As such, for simplify the mindset, each time we focus on implementing one operator for WebGPU and validate the operator implementation using the automated tests. Furthermore, for correctness, we also want to make sure after putting the newly implemented operator everything still works as well as previously. Therefore, in Section III-D, we maintain an initial CUDA compute graph whose correpsonding operators are gradually offloaded to WebGPU once a new WebGPU operator is implemented, and thus we perform model-wise validation in each step. The gradual offloading mechanism allows developers to focus on a minimal unit (i.e., one operator) each time, which minimizes the debugging scope, and once it passes model-wise validation we are confident that it will not bring hidden bugs to further steps. In Section III-E, TapML also simplifies these procedures through its universal runtime for flexible cross-platform interaction. The universal tooling stack in TapML can easily run remote computation as if locally and exchange useful information (e.g., debugging message) with the target platform.
III-B Prerequisites
While TapML is general, it makes the following lightweight assumptions to the ML framework using TapML:
Compute graph abstraction: TapML requires the developing system to model ML computation using a compute graph, i.e., a directed graph of tensor operators, each of which transforms a set of input tensors to output tensors. The abstraction of the compute graph allows for straightforward instrumentation of input and output tensors during test carving (in Section III-C). Viewing an operator as a minimal computation unit also offers a decent granularity for implementation, validation, and offloading to the intended platform (in Section III-D). This assumption is reasonable as the concept of compute graphs has long existed in various major ML frameworks and compilers, such as TensorFlow [23] and PyTorch [5].
Reference implementation: TapML for test automation depends on a mature platform as a reference to cross-check the implementation of the target platform. For example, a mature platform can be the host environment (e.g., x86/x64 CPU) or the NVIDIA ecosystem (e.g., CUDA). These platforms have been engineered for decades, achieving the highest level of coverage and reliability. The assumption is also practical given that CPU and CUDA are the de facto platforms supported by major ML frameworks since the early days.
III-C Test Carving
Figure 5: Carving intermediate inputs and outputs to a test case.
Test Carving [14] is a technique to automatically extract fine-grained unit tests from a given system-level test. Such unit tests are used to replay the intermediate behaviors (e.g., individual function invocations) that happened during the system test. To achieve this, carving techniques instrument [24] a system test to record invocations of desired functions along with their inputs and outputs. By knowing the function, inputs (e.g., arguments and context), and outputs (i.e., test oracle of differential testing [15]), a test carver synthesizes a unit test to reproduce the intermediate step of the system test.
In model deployment, a system test refers to the end-to-end validation of a full model, where the model is given some semantically meaningful inputs (e.g., pictures and sentences) and is expected to produce reasonable end-to-end outputs (e.g., correctly completing a motto). From the high level, TapML carves operator-level unit tests by instrumenting the end-to-end model execution on the source platform, to obtain reliable test inputs and reference outputs. Specifically, Figure 5 illustrates the carving process. First, the original compute graph on the source platform is modified such that each operator $\mathcal{F}_{S}$ is wrapped by an instrumentor, highlighted in the blue-green box. The instrumentation includes two instructions: (i) a pre-operator instruction to collect the operator call $\mathcal{F}$ and runtime inputs $\mathcal{I}$ feed to the instrumented operator; and (ii) a post-operator instruction to document the produced output $\mathcal{O}=\mathcal{F}_{S}(\mathcal{I})$ as the oracle from the source platform. Hence, each operator per execution forms a unit test as $t=(\mathcal{F},\mathcal{I},\mathcal{O})$ which can be replayed to validate the target-platform implementation. Such a source-to-target cross-checking mechanism also indirectly implements Differential Testing [15].
Notably, the model-wise inputs for bootstrapping the instrumented system test should be semantically meaningful, e.g., real-world data from the evaluation test set. This is important as it allows developers to focus the precision regression over realistic inputs, whereas synthetic test inputs used in the bottom-up pipeline is inadequate to validate the numeric precision stability due to the inconsistent distribution to natural inputs, though convenient as it is challenging to specify meaningful high-dimensional embeddings as intermediate inputs.
III-D Gradual Target Offloading
1
2
3Function GradualTargetOffloading($\mathcal{M}_{S}$):
4       $\langle\mathbb{F}_{S},\mathcal{G}\rangle\leftarrow\mathcal{M}_{S}$
5       $\mathbb{F}_{T}\leftarrow\{\}$
6       $\mathbb{T}=\{\mathcal{F}_{i}\mapsto\mathcal{T}_{i};i\in 1\cdots\lVert\mathbb{F}_{S}\rVert\}\leftarrow$Carving($\mathcal{M}_{S}$)
7       for $\mathcal{F}_{Si}\in\mathbb{F}_{S}$ do
8             $\mathcal{F}_{Ti}\leftarrow$ Initial implementation of $\mathcal{F}_{i}$
9             $\mathcal{T}_{i}\leftarrow\mathbb{T}[\mathcal{F}_{i}]$
10             while OpWiseValidation($\mathcal{T}_{i}$, $\mathcal{F}_{Ti}$) do
11                   Debugging $\mathcal{F}_{Ti}$
12
13            $\mathbb{F}^{*}\leftarrow$ $\mathcal{F}_{Sk}\mapsto\mathcal{F}_{Tk}(\mathbb{F}_{S});k\in\arg\left\{\mathbb{F}_{T}\cup\{\mathcal{F}_{Ti}\}\right\}$
14             while ModelWiseValidation($\langle\mathbb{F}^{*},\mathcal{G}\rangle$) do
15                   Debugging $\mathcal{F}_{Ti}$
16
17            $\mathbb{F}_{T}\leftarrow\mathbb{F}_{T}\cup\{\mathcal{F}_{Ti}\}$
18
19      $\mathcal{M}_{T}\leftarrow\langle\mathbb{F}_{T},\mathcal{G}\rangle$
20       return $\mathcal{M}_{T}$
21
22
23Function OpWiseValidation($\mathcal{T}$, $\mathcal{F}$):
24       for $t_{i}\in\mathcal{T}$ do
25             $\langle\mathcal{I},\mathcal{O}\rangle\leftarrow t_{i}$
26             if $\mathcal{F}(\mathcal{I})\neq\mathcal{O}$ then
27                   return False
28
29
30      return True
31
Algorithm 1 Gradual Target Offloading Procedure
TapML employs the Gradual Target Offloading mechanism to progressively transfer the compute graph from a standalone host to a standalone target platform. Algorithm 1 details steps of the Gradual Target Offloading procedure. In the high level, Gradual Target Offloading takes a source-platform compute graph $\mathcal{M}_{S}$ as input and migrates it to a target-platform compute graph $\mathcal{M}_{T}$. Recall Section III-B that $\mathcal{M}_{S}$ is a reference implementation from a mature source platform, assumed to be available beforehand. Both $\mathcal{M}_{S}$ and $\mathcal{M}_{T}$ can be initially represented using a general platform-agnostic representation shown in Figure 4 but they are specialized to different code targets. Specifically, $\mathcal{M}_{S}$ can be defined by a list of operators $\mathbb{F}_{S}$ and their topology information $\mathcal{G}$ (Algorithm 1). Therefore, to obtain $\mathcal{M}_{T}$ the goal is to transform $\mathbb{F}_{S}$ to $\mathbb{F}_{T}$ starting from an empty set (Algorithm 1). To start with, we perform test carving (Section III-C) to get a test suite where each operator maps to its corresponding unit tests (Algorithm 1). For each operator $\mathcal{F}_{i}$ in $\mathbb{F}_{S}$ we run following migration steps: (i) We draft an initial implementation for $\mathcal{F}_{i}$ on the target platform, i.e., $\mathcal{F}_{Ti}$ (Algorithm 1); (ii) we test and debug the implementation of $\mathcal{F}_{Ti}$ using the carved tests until the operator-wise validation passes (Algorithms 1 and 1); (iii) after operator validation, we perform model-wise validation over a hybrid compute graph $\langle\mathbb{F}^{*},\mathcal{G}\rangle$, by substituting corresponding parts in $\mathbb{F}_{S}$ with implemented target-platform operators, i.e., $\left\{\mathbb{F}_{T}\cup\{\mathcal{F}_{Ti}\}\right\}$ (Algorithms 1, 1 and 1); and (iv) lastly we append the tested $\mathcal{F}_{Ti}$ to $\mathcal{M}_{T}$ to finalize the migration of $\mathcal{F}_{i}$ (Algorithm 1). As such, once all target-platform operators are implemented and tested (i.e., $\mathbb{F}_{T}$), we wrap up the procedure by returning a compute graph $\mathcal{M}_{T}$ fully supported on the target platform (Algorithms 1 and 1).
Figure 6: Sample validation code using TapML APIs.
III-E Universal Runtime
As is shown in Figure 4, we need a runtime system to transfer data (e.g., device code and test inputs) and perform differential testing. To this end, we build a universal runtime to unify these procedures to facilitate cross-platform data transfer and operation. We argue that it is crucial to unify related behaviors through a standardized interface, which provides a reliable programming model and maximizes code reuses.
Exemplification. Figure 6 demonstrates the interface via an example of differential testing on CUDA and WebGPU implementations. Specifically, the tester takes a transformer_block operator module and a set of computational inputs as the function input (Line 3), based on which we compile the module to both CUDA and WebGPU platforms and perform differential testing (Line 21). In Line 4, the local function run compiles the transformer_block module on a given device, returning the results of running the compiled module over the given inputs. More specifically, we first compile the target device code locally (Line 6) and upload the compiled device code to the remote device via tapml.upload (Line 7), which returns a handle (i.e., reference) to the remote library. Next, we convert and transfer the test inputs from the local host to the device platform as tml_inputs in Line 10. To obtain the output on the device, we also need to allocate space for the output tensor tml_output in Line 12. As such, in Line 15 and 16 we call the remote device function which transforms tml_inputs to tml_output, which are transferred back to the local host in Line 19 to perform concrete testing.
Implementation. The core of the runtime is communication, i.e., how to transfer data and instructions across multiple platforms. Consequently, we develop a lightweight Remote Procedure Call (RPC) framework to manage host-device data transfer and perform device function calling. Additionally, error handling is a critical yet challenging aspect of universal runtime systems. When an error occurs, the runtime is designed to collect error information as much as possible, with complete trackback information, and transmit these debugging messages back to the host before device termination.
IV Evaluation
We evaluate TapML via the following research questions:
  • RQ1 (Section IV-B): How is the overall model support and timeline by using TapML?
  • RQ2 (Section IV-C): How is the performance and overhead of using the universal runtime component?
  • RQ3 (Section IV-D): What are the bugs and challenges when developing emerging ML systems?
  • RQ4 (Section IV-E): What are the lessons and best practices of using TapML?
IV-A Setup
{tblr}
row2,4,5,8,9=bg=azure9, colspec=Q[c,white] — Q[c] Q[c] Platform Hardware
[…]
WebGPU Apple Silicon M1 Pro
(MacBook Pro)
[…]
(Xiaomi 14 Pro)
[…]
(iPhone 14 Pro)
ROCm Radeon 7900 XTX
TABLE I: Mature and emerging platforms in TapML.
Platforms. To demonstrate the versatility and adaptability of TapML across a diverse range of configurations, we evaluate TapML over platforms shown in Section IV-A. The corresponding hardware selection spans from a GPU-intensive workstation to energy-efficient mobile devices. Furthermore, we consider a rich set of mature and emerging runtime platforms below:
  1. 1.
    CUDA [25]: A parallel computing platform and application programming interface model for NVIDIA GPUs.
  2. 2.
    ROCm [26]: An open-source platform by AMD for GPU-accelerated computing, analogous to NVIDIA’s CUDA but for AMD GPUs. It provides a comprehensive suite of tools and libraries for high-performance, parallel computing applications across various domains.
  3. 3.
    Metal [11]: Apple’s low-level API for optimizing graphics and computing on Apple ecosystems such as iOS and macOS, offering high performance and efficiency.
  4. 4.
    WebGPU [12]: A web standard providing cross-platform, low-level graphics and computation APIs and capabilities on web browsers.
  5. 5.
    Vulkan [27]: A cross-platform graphics and compute API, providing high-efficiency, low-overhead access to modern GPUs on various devices.
  6. 6.
    OpenCL [28]: An open, royalty-free standard for cross-platform, parallel programming of diverse processors found in personal computers, servers, mobile devices, and embedded platforms.
Base framework. We experiment TapML based on Apache TVM [6], an open-source ML compiler stack with robust support for CPU and NVIDIA GPU platforms. Specifically, we build a universal ML inference engine targeting emerging models and platforms as part of the MLC-LLM project at https://github.com/mlc-ai/mlc-llm on top of TVM.
IV-B RQ1: Support Overview
{talltblr}
[…]
CUDA Vulkan ROCm Metal WebGPU OpenCL Metal
[…]
StarCoder-1B Apr 24 2023 June 23 2023 430.2 157.8 388.9 86.0 49.1 23.1 38.1
RedPajama-3B May 5 2023 May 19 2023 283.0 131.2 245.1 47.5 29.7 14.1 19.5
InternLM-7B July 6 2023 Feb 28 2024 138.8 96.3 148.4 26.0 15.7 8.3 3.2
WizardMath-7B Aug 11 2023 Aug 12 2023 147.2 106.4 168.6 27.4 17.2 8.8 5.4
Baichuan2-7B Sept 6 2023 Feb 16 2024 129.1 97.1 153.5 26.9 14.1 10.8 1.9
TinyLlama-1.1B Sept 16 2023 Sept 21 2023 577.9 279.2 487.5 103.3 63.1 29.9 42.3
Mistral-7B Sept 27 2023 Sept 28 2023 128.1 85.1 139.5 24.7 15.5 6.3 4.1
StableLM-3B Sept 29 2023 Oct 3 2023 281.2 135.8 240.8 47.4 34.2 19.9 19.6
[…]
Qwen-1.8B Nov 30 2023 Jan 2 2024 306.1 169.0 282.6 74.4 28.3 13.7 32.1
Phi-2 Dec 12 2023 Dec 28 2023 171.0 107.6 194.7 36.3 23.5 11.2 14.4
Qwen1.5-4B Feb 5 2024 Feb 13 2024 179.1 102.9 167.0 38.1 19.2 10.2 15.9
Orion-14B Jan 20 2024 Mar 4 2024 71.8 41.9 77.6 OOM OOM OOM OOM
Gemma-2B Feb 21 2024 Feb 21 2024 266.0 127.8 201.4 50.1 23.2 14.6 22.0
[…]
In a year-long effort (starting from April 2023), we based on TapML have deployed 82 models (i.e., validated) of 17 model architectures over 5 emerging platforms. To our knowledge, to date TapML offers the most comprehensive range of platform support. More importantly, TapML also becomes the fastest and even sole supporter on a line of important scenarios:
  1. 1.
    The first LLM on mobile GPUs: Thanks to TapML, we are the first in open source to successfully deploy LLMs on mobile GPUs of Android and iOS devices. By maximizing the efficiency of mobile GPUs via Metal and OpenCL, running large LLMs on mobile phones becomes a possible and decent experience, e.g., running Baichuan2-7B on Snapdragon 8 Gen 3 at over 10 tokens per second. Our pioneering work motivates the community to develop native and mobile LLM applications without the fear of sacrificing user privacy.
  2. 2.
    The first LLM on WebGPU: TapML allows us to be the first in public to run LLMs on WebGPU, despite its status as an experimental feature in Chrome at the time of development. Enabling emerging ML on WebGPU is of significant broader impact as it allows everyone to run powerful models as easily as opening a web page. Notably, nearly one year after our effort, Google introduced the MediaPipe [29] via LLM Inference API on TensorFlow.js [16] for WebGPU and mobile platforms. We believe TapML approach can complement and speed up the development of these related efforts.
  3. 3.
    The only Orange Pi LLM: To date, we are the only team that has successfully deployed LLMs on the Orange Pi 5 [30], an edge device and embedded system in \$100. Yet, we achieve a practical operational speed of 2.5 tokens per second for running Llama2-7B and 5 tokens per second for RedPajama-3B using the OpenCL APIs to maximize the edge GPU on Orange Pi. This inspires the community to develop affordable LLM applications on edge scenarios.
Thanks to the techniques and design choices in TapML, these highlights are achieved by our small team. Through test carving TapML frees developers from writing tedious test cases. Meanwhile, the gradual offloading mechanism in TapML minimizes the mental burden of developers by allowing them to focus on the minimal implementation unit of single operators each time until it is fully validated. Lastly, the universal runtime becomes a huge time saver since interacting with such emerging and under-developed platforms is made as easy as calling Python functions locally.
More specifically, LABEL:tab:model_support details the availability and performance of a selection of models and platform supports. Thanks to TapML, we quickly deploy recent and popular models within a short period, e.g., up-to same-day support, demonstrating the effectiveness of TapML in optimizing development productivity. More importantly, all of the deployed models are extensively validated by carved tests from TapML as well as common end-to-end tests and achieve outstanding processing performance to serve these models efficiently.
IV-C RQ2: Overhead of TapML
{tblr}
row2,4,6,8=bg=azure9, colspec=Q[c,white] — Q[l] Q[r] Q[c] Q[r] Development step Platform Cost O(1)?
[…]
Weight loading WebGPU 1.6s ✗
End-to-end validation WebGPU 4.6s ✗
[…]
Weight loading CUDA 1.6s ✓
Test carving CUDA 5.2s ✓
Operator validation WebGPU 0.1s ✗
TABLE II: Testing time breakdown in TapML exemplified by deploying Llama2-7B on WebGPU with CUDA reference. “O(1)” stands for one-time effort.
In this section, we systematically examine the overhead incurred by TapML during development. For clarity, we exemplify the overhead analysis via a case study of deploying Llama2-7B on WebGPU by using CUDA as the reference platform, running on a workstation with an AMD Ryzen 9 5900X CPU and an NVIDIA RTX 3080 GPU. Specifically, Llama2-7B includes 32 decoder blocks and when being compiled to the general representation of our base framework, it leads to 456 operator kernels in 18 operator types. Consequently, to support Llama2-7B, we need to implement test 18 operator kernel functions and pass $456\times k$ operator-level tests, where $k$ is determined by the number and sizes of sequences to be completed.
Section IV-C depicts the testing time breakdown of supporting Llama-2 given 6 prefilling token and 1 decoding token. Specifically, testing the model includes two types of costs:
  1. 1.
    Rooting time: The time spent on running a model-wise validation on the target platform (i.e., WebGPU), including compiling the model into WebGPU code, initializing the model on WebGPU, and running the model on WebGPU with model-wise test inputs. These steps are used for end-to-end model validation which is indispensable regardless of the development methods employed. Specifically, the bottleneck comes from model compilation where compiling Llama2-7B fully on WebGPU results in 87% of the end-to-end testing time.
  2. 2.
    Overhead time: Besides routine end-to-end validation, we need to obtain a full CUDA compute graph for test carving and gradual offloading. While compiling, loading, and test-carving Llama2-7B on CUDA can take about 1 minute, these steps are one-time efforts and can facilitate the testing of all operators. For operator-wise validation, we need to use the universal runtime to run carved tests for each operator, whose cost is tiny, e.g., on average 0.1 seconds per operator.
To conclude, the overhead introduced by TapML is negligible compared to the significant improvement of productivity when developing emerging ML systems.
IV-D RQ3: Bugs in Emerging ML Systems
TapML aims to facilitate the model development on various emerging platforms. These next-gen platforms are so new that oftentimes they come with limited ecosystem and test coverage, leading to challenging bugs and sophisticated debugging experience. Consequently, in addition to our technical contribution, we share our concrete experiences of debugging two types of bugs encountered during our development.
Bugs by platform runtime. The emergence of new computing platforms often brings with it a degree of instability, including inconsistent or incomplete implementations to its specifications. For instance, we found one of our WebGPU programs runs flawlessly on Google Chrome but produces wrong results on the Technology Preview (Release 188) version of Safari on the same MacBook. Our diagnosis shows that such discrepancies can be attributed to the incomplete WebGPU API support on Safari, such as onSubmittedWorkDone, whose incompleteness leads to a synchronization bug in which unready buffers were accessed. Furthermore, the base framework by default reuses shared memory buffers of different data types, which is okay for CUDA but introduced Vulkan compilation failures as this feature is not permitted by Vulkan’s design. Additionally, WebGPU is subject to varying constraints across different devices. For example, when deploying Gemma on WebGPU of the MacOS system, we assumed the maximum number of threads to be the same as that of the native Metal runtime. However, this leads to GPUPipelineError because WebGPU at most allows for 256 threads on Apple Silicon which is much smaller than our assumed value. Similarly, on the PC end, WebGPU allows for 1024MB maximum storage buffer sizes, yet it is limited to 128MB for mobile devices that over-requesting buffer sizes leads to an error of “Invalid BindGroup”.
Bugs by base framework. Our base framework, i.e., TVM, is a compilation-based framework that compiles an ML model from its internal representation into target code. When extending our base framework, a range of issues surfaced, predominantly from compiler errors from both the base and extended parts of the framework. For instance, we observed that our generated Metal code does not perform checks for unaligned data, which leads to incorrect results due to accessing unknown data. Meanwhile, missing instruction support in code generation is a major error type discovered by our carved tests. For example, certain compilation failures were encountered when offloading a matrix multiplication operator to the Metal platform, because of unsupported SIMD instructions such as the selection instructions. Furthermore, we also encountered performance issues. For example, our Metal code-generation pipeline initially did not support warp-level instructions, such as simd_shuffle, and therefore vanilla instructions are used which leads to sub-optimal performance. Meanwhile, we also used to assume an inefficient setting of warp sizes on Apple Silicon devices, leading to undesired processing speed. All of these errors are quickly located thanks to the test carving and gradual offloading mechanism in TapML, which isolates the debugging scope to a single operator each time.
Besides errors from system implementations, the rest of the issues pertain to numerical discrepancies that arise during runtime, making them considerably more challenging to pinpoint. For example, we detected numerical errors when running LLMs with int8 quantization on the Metal platform, yet the same quantization method works perfectly on other platforms. To isolate this issue, we leveraged automated unit tests carved from the CUDA implementation of the operators and compared the results with those from the Metal platform. As such, we quickly identified the root cause in the dequantization kernel, where an unexpected conversion from uint8x4 to uint32 was erroneously performed prior to converting to halfx4. The correct procedure should have been a direct conversion from uint8x4 to halfx4. Remarkably, this debugging process was completed within just two hours, showcasing the debuggability of TapML.
The presented cases empirically demonstrated the huge inconsistency and complexity of various platforms, posing significant challenges in model deployment. Nevertheless, the adoption of a top-down methodology has significantly enhanced our ability to support emerging models progressively. Initially, our adaptation process could span up to a few months based on a traditional bottom-up flow; however, with refinements to our approach, we have reduced this time to even a single day, regardless of the additional development complexity involved by the new runtime environments or backend systems. This agile response capability underscores the effectiveness of the top-down approach in the dynamic landscape of language model development.
IV-E RQ4: Lessons and Best Practices
We present a practical exploration of real-world debugging scenarios, illustrating how the integration of TapML with a top-down methodology can be effectively applied to complex, real-life cases. We showcase two best practices of using TapML, corresponding to our two main sub-techniques.
Best practice of test carving: Debugging runtime numerical errors, particularly those that occur sporadically, presents a significant challenge. Such errors were encountered when developing the Phi-2 model on Metal, leading to the emergence of NaN outputs. Interestingly, this problem did not happen on CUDA and was only induced when using certain prompts.
To address this, we employed top-down debugging via TapML. By using our test carvers to automatically break down the CUDA Phi-2 model into individual reliable tests, we automatically examine the correctness of the Metal implementation at a granularity of operators, which allowed us to quickly pinpoint the GeLU operator as the source of erroneous output. Further investigation revealed that the root cause was the Metal platform’s numerically unstable fast-math implementation of the tanh function, which could yield NaN values for inputs exceeding 45. We resolved the issue by disabling the fast-math option on the Metal platform, which is replaced by a numerically stable version of the tanh reimplemented by us.
In contrast, a bottom-up strategy for debugging would likely have entailed writing unit tests for each operator, typically with random inputs. Such a method, however, would not always trigger the specific error we encountered as it depends on particular activation values that can be induced by real-life input. Meanwhile, it is also hard to pin down the error in end-to-end integration. Our experience underscores the importance of a targeted, top-down method to isolate and resolve elusive numerical errors in complex computational models.
Best practice of gradual offloading: RWKV [31] is an innovative neural network architecture with RNN architecture and Transformer-level performance. Distinct from the prevailing Transformer architectures, RWKV introduces a unique and specialized scan kernel, namely the wkv kernel. This kernel does not lend itself to straightforward representation as a mere aggregation of conventional operators, which requires support from scratch to deploy on nascent platforms.
The gradual target offloading mechanism in TapML facilitates the deployment by allowing us to focus on individual operators. Initially, we implemented the wkv kernel by describing its IR in our base framework, which is then offloaded to the target platform, with every else model component maintained on the CUDA platform. As such, we isolate the newly implemented wkv target kernel into the CUDA graph and through end-to-end model validation can fully validate if the new support can tackle real-life inputs. Thanks to the universal runtime, the detailed steps such as data transfer and remote code execution are all done in a Python environment as if locally, largely lowering the programming barrier. Consequently, by using TapML, we become the first and sole team to run RWKV models across a wide range of platforms.
In contrast, a bottom-up scheme would have necessitated the completion of the entire RWKV model before end-to-end testing on the target platform. This could lead to significant challenges in troubleshooting if the eventual integration turns out to be erroneous. Furthermore, the wkv kernel’s dependency on real-world inputs and the state of preceding tokens complicates the creation of unit tests. Given these intricacies, comprehensive end-to-end validation or automatic tests emerge as viable options for progressive model deployment.
V Related Work
V-A Emerging ML Models and Platforms
Machine Learning has demonstrated remarkable capabilities across a variety of tasks. Specifically, Large Language Models [1, 32, 33, 34] have shown proficiency in linguistic tasks by accurately predicting subsequent tokens. Meanwhile, diffusion models [35, 36, 37, 38, 39] are adept at producing realistic images and videos in response to textual prompts. However, deploying these models presents significant challenges. They not only require substantial computational resources but also necessitate repeated executions within a single request.
As these models increase in size and complexity, the desire to deploy them broadly conveniently, and powerfully persists. Major high-performance-computing corporations such as NVIDIA, AMD, Intel, and Google have released cutting-edge AI accelerators. Concurrently, startups like SambaNova, Cerebras, and Tenstorrent are introducing chips with innovative architectures. Additionally, mobile SoCs have incorporated specialized Neural Processing Units (NPUs), e.g., Apple’s Neural Engine and Qualcomm’s Hexagon Tensor Processor.
To close the gap between hardware performance and programmability, open APIs such as Vulkan and OpenCL enable compatibility with different GPU types, allowing for the construction of universally deployable models on supported hardware without complicated dependency chains. More recently, WebGPU unlocks the GPU power within web browsers. While these runtimes offer considerable convenience in production, they also pose development challenges due to the nascent state of their associated development and debugging tools.
At the higher level, several recent frameworks have been developed to catch up with the rapid evolution of models. Examples include vLLM [40], Hugging Face Transformers [41], llama.cpp [42], TensorRT-LLM [43], and MediaPipe [29]. However, these frameworks often focus on major or specific platforms, e.g., CUDA for NVIDIA GPUs. TapML aims to close the gap by pioneering the support and optimization of underrepresented emerging platforms, to maximize the possible scenarios of ML applications. Meanwhile, instead of cherry-picking specific platforms, the top-down methodology in TapML unleashes the productivity of developing emerging ML systems, which allows us to have the most comprehensive yet timely support of emerging platforms in public.
V-B Testing and Debugging in ML Frameworks
Software testing techniques for ML frameworks and compilers have been well-studied in recent years to improve their reliability. Specifically, due to the huge complexity of large-scale systems, fuzzing [44] has been applied to generate random ML models as test programs, to exercise the compilation and execution pipeline of ML frameworks. To this end, recent work has been focusing on generating valid and diverse ML models [45, 46, 47, 48, 49, 17, 18] (e.g., PyTorch programs) or IRs [50, 51]. For example, NNSmith [17] applies a solver-aided approach to define operator constraints and generate valid graphs through SMT solving. Besides test model generation, recent work also improves the reliability of ML systems through translation validation [52] and discovering more oracles [53, 54, 55]. Rather than generating random models to test the whole system comprehensively, TapML focuses on testing a vertical path to deploying one model each time.
Debugging techniques have been used in ML system development to accelerate troubleshooting. PolyGraphy [22] is a tool aiming to diagnose a small error-inducing subgraph out of an ONNX model [56] that fails on NVIDIA TensorRT [4], via test input reduction [57]. Specifically, PolyGraphy either performs bisection or linearly removes the top and bottom layers to get smaller sequence models to reproduce the error. However, such a test input reduction mechanism is neither optimal nor efficient, as it assumes the input model to be a sequence that can actually be a graph and requires significant attempts of recompilation and testing. TapML addresses the limitations by opening the black box of the DL framework to directly isolate errors in small minimal subgraphs through test carving [14].
VI Threats to Validity
External. The main concern of external validity is if TapML can be applied to develop emerging ML systems. To alleviate the concern, we built TapML and conducted the experiments by diversifying the experimental configurations, including the supported models, platforms, and involved developers. Specifically, we have successfully applied TapML to deploy 82 important models on 5 emerging and serious platforms successfully. Meanwhile, to our best effort, we employ as many as 11 developers involved in the development and study. The employed developers have different years of experience ranging from 0 to over 5 years and we have been practicing the methodology for over a year.
Internal. A potential threat to internal validity is that there may be faults in our implementation and experiments. We perform systematic unit testing and code reviews during our development to alleviate this concern.
VII Conclusion
TapML aims to address the critical software development challenges posed by the rapid evolution of ML models and the emergence of new computing platforms. By adopting a top-down approach and utilizing the universal runtime, TapML streamlines the deployment process on diverse platforms, significantly enhancing developer productivity. This method, characterized by automated testing and a migration-based strategy, has proven to be effective in real-world applications, as evidenced by the successful deployment of 82 emerging models across 5 novel platforms. The comprehensive case studies derived from these deployments offer great insights and best practices for developing ML systems on emerging platforms. Ultimately, we show that the TapML approach accelerates the deployment process while ensuring the reliability and efficiency of emerging models, bridging the gap left by traditional methods and existing frameworks.